solve this quadratic equation 6x2-7x-3=0
Answer by nerdybill(7384)

6x^2-7x-3=0
can be factored:
6x^2+2x-9x-3=0
(6x^2+2x)-(9x+3)=0
2x(3x+1)-3(3x+1)=0
(3x+1)(2x-3)=0
x = {-1/3, 3/2}